# X-Vine Models for Multivariate Extremes

### v1

@@ 7.3 Model selection for X-vines
Model selection for X-vines based on a random sample $\bm{X}_{1},\ldots,\bm{X}_{n}$ from $F$ as in Section 7.2 involves two steps:
@@
Selecting parametric (tail) copula families given a truncated regular vine sequence
We first consider the specification of $\mathcal{R}$ in Definition 5.6 in tree $T_{1}$. Let $\mathcal{B}_{\mathcal{R},1:T}=(\mathcal{B}_{\mathcal{R},1},\ldots,\mathcal{B}_{\mathcal{R},T})$ be the list of candidate bivariate tail copula families. For each edge $e=(e_{a},e_{b})\in E_{1}$, specifying $\mathcal{R}$ involves choosing the bivariate tail copula family among $\mathcal{B}_{\mathcal{R},1:T}$. Similar to the idea of averaged maximum pseudo-likelihood estimates in Section 7.2, we use the averaged-AIC value for selecting bivariate tail copula families. For each edge $e$ and each $\mathcal{B}_{\mathcal{R},t}$, $t=1,\ldots,T$, we obtain two maximum pseudo-likelihood estimates $\hat{\theta}_{e}^{(t,a_{e})}$ and $\hat{\theta}_{e}^{(t,b_{e})}$, derived through the maximisation of log-likelihood functions on product spaces; see (7.4). The averaged-AIC value is
$\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{R},t}}\right)=2\nu^{(t)}-\tfrac{1}{2}\left\{\log\mathcal{L}_{\mathcal{R},e}\left({\hat{\theta}_{e}^{(t,a_{e})}}\right)+\log\mathcal{L}_{\mathcal{R},e}\left({\hat{\theta}_{e}^{(t,b_{e})}}\right)\right\},$
where $\nu^{(t)}$ is the number of parameters in $\mathcal{B}_{\mathcal{R},t}$. We select the bivariate tail copula family with the lowest averaged-AIC among $\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{R},1}}\right),\ldots,\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{R},T}}\right)$.
Similarly, we specify $\mathcal{C}$ in Definition 5.6 in $T_{j}$ for $j\in\left\{{2,\ldots,d-1}\right\}$. Let $\mathcal{B}_{\mathcal{C},1:S}=(\mathcal{B}_{\mathcal{C},1},\ldots,\mathcal{B}_{\mathcal{C},S})$ denote the list of candidate bivariate parametric copula families, including the independence copula (without parameters). Specifying $\mathcal{C}$ involves selecting the bivariate parametric copula family among $\mathcal{B}_{\mathcal{C},1:S}$ for each edge $e=(e_{a},e_{b};D_{e})\in\bigcup_{j=2}^{d-1}E_{j}$. We do so by maximising the AIC, following the common practice in vine copula modelling; see Brechmann (2010) and (Czado, 2019, Section 8.1). For each edge $e$, we estimate the associated parameter (vector) $\theta_{e}^{(s)}$ by maximising the log-likelihood (7.5), and we calculate the corresponding AIC over the list of the bivariate parametric copula families $\mathcal{B}_{\mathcal{C},s}$, $s=1,\ldots,S$,
$\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{C},s}}\right)=2\nu^{(s)}-2\log\mathcal{L}_{\mathcal{C},e}\left({\hat{\theta}_{e}^{(s)}}\right),$
where $\nu^{(s)}$ is the number of parameters in $\mathcal{B}_{\mathcal{C},s}$, and $\nu^{(s)}=0$ if $\mathcal{B}_{\mathcal{C},s}$ contains only the independence copula. We then choose the bivariate parametric copula family with the lowest value among $\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{C},1}}\right),\ldots,\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{C},S}}\right)$.
@@
Selecting the regular vine sequence
@@
To select the first tree, $T_{1}=(N_{1},E_{1})$, on the node set $N_{1}=[d]$, we follow a procedure as in Engelke and Volgushev (2022) and Hu et al. (2023). For every pair $\left\{{a,b}\right\}$ of distinct elements in $[d]$, let $w_{a,b}$ be a nonnegative weight derived from the data. We will be using an empirical version of the tail dependence coefficient $\chi_{a,b}=R_{a,b}(1,1)$, namely we will set $w_{a,b}$ equal to
$\widehat{\chi}_{a,b}=\frac{1}{k}\sum_{i=1}^{n}\operatorname{\mathds{1}}\left\{{\widehat{U}_{i,a}\leqslant k/n,\widehat{U}_{i,b}\leqslant k/n}\right\},$
where $\widehat{U}_{i,a}$ and $\widehat{U}_{i,b}$ are defined as in (7.2). Another possible edge weight would be the empirical extremal variogram as in Engelke and Volgushev (2022). The first tree is then selected as the maximum-spanning tree based on the edge weights $w_{a,b}$, that is,
$T_{1}=\operatorname*{arg\,max}_{\text{spanning tree }T=(N_{1},E)}\sum_{e=\left\{{a,b}\right\}\in E}w_{a,b}.$
For each edge $e$ in $T_{1}$, we then select parametric tail copula densities and estimate their parameters as previously explained.
Next we describe, for $j\in\left\{{2,\ldots,d-1}\right\}$, how to select the tree $T_{j}=(N_{j},E_{j})$ given the already selected trees $T_{1},\ldots,T_{j-1}$ and the parametric families and parameter estimates for edges $e\in E_{1}\cup\cdots\cup E_{j-1}$. The edge set of $T_{j}$ is necessarily $N_{j}=E_{j-1}$. Let $E_{j,P}$ be the set of pairs of elements of $N_{j}$ that satisfy the proximity condition (Definition 2.2); that is, every element $e$ of $E_{j,P}$ is a pair $\left\{{f,g}\right\}$ of distinct elements in $N_{j}$ so that $f\cap g$ is a singleton. For such $e\in E_{j,P}$, we can then consider the conditioning set $D_{e}$ and the conditioned set $\mathcal{C}_{e}=\left\{{a_{e},b_{e}}\right\}$ as before and, by recursion, compute $\widehat{U}_{i,a_{e};D_{e}}$ and $\widehat{U}_{i,b_{e};D_{e}}$ for $i\in N_{D_{e}}$ as in (7.6). Based on these, we compute an edge weight $w_{e}$; we will take the absolute value of the empirical Kendall’s tau, $|\widehat{\tau}_{e}|$. We then select $T_{j}$ as the maximum spanning tree on the node set $N_{j}$ with edges restricted to $E_{j,P}$, that is,
$T_{j}=\operatorname*{arg\,max}_{\text{spanning tree $T=(N_{j},E)$ with $E\subseteq E_{j,P}$}}\sum_{e\in E}w_{e}.$
For each edge $e\in E_{j}$, we further select the parametric copula families and estimate their parameters as before.
Selecting truncated regular vine tree sequences
The model selection procedure described above may set all pair-copulas to the independence copula in the subsequent trees from $T_{q+1}$ to $T_{d-1}$ for a truncation level $q\in\left\{{1,\ldots,d-2}\right\}$. In this case, the resulting model corresponds to the truncated X-vine model in Definition 5.10. Besides an information criterion, we consider two additional criteria for selecting the independence copula at an edge $e$: when the effective sample size $|K_{D_{e}}|=n_{D_{e}}$ falls below a certain low value, and when the absolute value of the empirical Kendall’s tau is close to zero. Sparsity, induced by a smaller effective sample size, is more likely when the total sample size $n$ is relatively smaller than the dimension $d$; recall that we condition on $(j-1)$ variables being extreme in tree $T_{j}$ as explained in Section 7.2.
@@
$\operatorname{mBIC}(q)=\sum_{j=2}^{q}\sum_{e\in E_{j}}\left[{\operatorname{\mathds{1}}(\theta_{e}\neq 0)\left\{{\log{n_{D_{e}}}-2\log\left({\frac{\psi_{0}^{j-1}}{1-\psi_{0}^{j-1}}}\right)}\right\}-2\log\mathcal{L}_{\mathcal{C},e}\left({\hat{\theta}_{e}}\right)-2\log\left({1-\psi_{0}^{j-1}}\right)}\right],$ (7.7)
with $\mathcal{L}_{\mathcal{C},e}$ as in (7.5). The selected truncation level $q^{*}$ is the value of $q$ in $\left\{{1,\ldots,d-1}\right\}$ that minimises $\operatorname{mBIC}(q)$. In practice, we will set $\psi_{0}=0.9$ as in Nagler et al. (2019).
@@ 8.1 Parameter estimation
We first consider the $5$-dimensional X-vine model $r$ in Fig. 2. The bivariate tail copula densities $r_{a_{e},b_{e}}$ in the first tree $T_{1}=(N_{1},E_{1})$ are chosen from the Hüsler–Reiss model, the negative logistic, logistic, and Dirichlet families (Section 4), while the bivariate copula densities $c_{a_{e},b_{e};D_{e}}$ in the subsequent trees $T_{j}=(N_{j},E_{j})$ for $j\in\left\{{2,3,4}\right\}$ are taken from the Clayton, Gumbel, and Gaussian copula families. We specify the parameter values for the bivariate parametric families as follows, including the tail dependence coefficient $\chi=R(1,1)$ or Kendall’s $\tau$ for comparison: $\theta_{12}=1.5$ $(\chi=0.54)$, $\theta_{23}=2$ $(\chi=0.71)$, $\theta_{24}=2.5$ $(\chi=0.68)$, $\theta_{45}=2$ $(\chi=0.62)$; $\theta_{13;2}=2$ $(\tau=0.50)$, $\theta_{34;2}=2.5$ $(\tau=0.60)$, $\theta_{25;4}=0.7$ $(\tau=0.49)$; $\theta_{14;23}=0.4$ $(\tau=0.17)$, $\theta_{35;24}=-0.3$ $(\tau=-0.19)$; and $\theta_{15;234}=0.1$ $(\tau=0.06)$. The formulas connecting the parameters of the families of tail copula densities with the value of $\chi$ are given in Appendix F.2.
@@
Figure 2: An X-vine specification $(\mathcal{V},\mathcal{R},\mathcal{C})$ on $d=5$ variables. The bivariate tail copula densities $r_{a_{e},b_{e}}$ in the first tree $T_{1}$ are selected from the Hüsler–Reiss (HR), negative logistic (NL), logistic (L), and Dirichlet (Diri) parametric families, each with specified tail dependence coefficients $\chi$. The bivariate copula densities $c_{a_{e},b_{e};D_{e}}$ in the subsequent trees $T_{2},T_{3},T_{4}$ are chosen from the Clayton (Clay), Gumbel (Gum), and Gaussian (Ga) copula families, each with specified Kendall’s tau $\tau$.
Implementing simulation algorithms as detailed in Section 6 and Appendix A, we generate multivariate inverted Pareto random samples $\bm{Z}_{1},\ldots,\bm{Z}_{n}$ associated with $r$ in Section 6. Rather than using these directly, we perform a rank transformation $\widehat{\bm{U}}_{1},\ldots,\widehat{\bm{U}}_{n}$ as defined in (7.1) with $\widehat{F}_{j}(Z_{i,j})=(\operatorname{rnk}_{i,j}-0.5)/n$, where $\operatorname{rnk}_{i,j}$ denotes the rank of $Z_{i,j}$ among $Z_{1,j},\ldots,Z_{n,j}$ for $j\in[d]$. We then take sub-samples $\{\widehat{\bm{Z}}_{i}\}_{i\in K}$ with $K=\bigcup_{j\in[d]}K_{j}$ and $K_{j}$ as in (7.3) to investigate simulation results with respect to threshold exceedances.
@@
Figure 3: (a) Box-plots ($\blacksquare$ for $T_{1}$blacksquare for $T_{2}$blacksquare for $T_{3}$blacksquare for $T_{4}$) of dependence measure estimates from specified bivariate parametric families for each edge across the four trees in the X-vine specification in Fig. 2. (b) Similar as the plot on the left, but now bivariate parametric families are selected sequentially from the data, yielding larger estimation uncertainty. The red lines indicate the specified parameter values.
@@ 8.2 Selecting parametric (tail) copula families
We assess algorithm effectiveness in selecting bivariate parametric families for each edge in each tree, using the same X-vine specification as in Fig. 2. While the flexibility of X-vine models allows us to consider any bivariate parametric family, we simplify the process by considering a catalogue of four candidate tail copula models for $T_{1}$: the Hüsler–Reiss, negative logistic, logistic model, and Dirichlet model, along with a catalogue of nine candidate pair-copula families for $T_{2},T_{3},T_{4}$: Independence, Gaussian, Clayton, Survival Clayton, Gumbel, Survival Gumbel, Frank, Joe, and Survival Joe copulas, as implemented in the VineCopula software package (Nagler et al., 2023).
@@
We also investigate how effective sample sizes change over tree levels shown in Table 0(b). We calculate averaged-effective sample size over 200 repetitions for each edge in each tree and express it as a percentage of the sample size $n$. The overall averaged-effective sample size decreases in the first level tree when considering stronger extremal structures (result is omitted). This reduction is due to the fact that, for a high threshold $u$, the probability $\operatorname{\mathsf{P}}[X_{a}>u\text{ or }X_{b}>u]$ decreases as the tail dependence between $X_{a}$ and $X_{b}$ increases. For the second level tree $T_{2}$, the averaged-effective sample is $k/n=5\%$, since the use of the rank transform implies (in the absence of ties) the identity $K_{j}=k$ for all $j\in[d]$.
@@
$T_{1}$ 7.30% 6.47% 6.60% 6.87%
$T_{2}$ 5.00% 5.00% 5.00%
$T_{3}$ 3.52% 3.38%
$T_{4}$ 3.12%
Table 1: (a) Proportion of correctly selected families for each edge in each tree. (b) Averaged-effective sample sizes as percentages for the same edges. The order of the components in Table 1, from left to right and then from top to bottom, corresponds to the x-axis order of box-plots in Fig. 3.
@@ 8.3 Tree selection and truncation
To explore truncated X-vine models, we use the mBIC with $\psi_{0}=0.9$ as in Nagler et al. (2019). For a specific iteration, we select the mBIC-optimal truncation level of $q^{*}=19$, indicated as a dotted line in Fig. LABEL:fig:mBIC50dim. We assess the goodness of fit by comparing pairwise $\chi$-values from the specified 50-dimensional X-vine model with those obtained from the fitted truncated X-vine model via Monte Carlo simulations in Fig. LABEL:fig:Chi50dim (for bivariate tail copula densities, closed-form expressions of $\chi$ are available for most well-known parametric tail copula density families; see Appendix F). In the case of the truncated X-vine model, the chi-plot in Fig. LABEL:fig:Chi50Trunc resembles that of the full model but exhibits more variability, particularly for lower $\chi$ values.
Figure 4: (a) The mBIC plotted across tree levels with a dotted line indicating the selected mBIC-optimal truncation level of $q^{*}=19$. (b) $\chi$-plot comparing pairwise $\chi$ values from the specified X-vine model to those obtained from the fitted X-vine model via Monte Carlo simulations. (c) Similar to (b), a $\chi$-plot for the truncated X-vine model.

### v2

@@ 7.3 Model selection for X-vines
Model selection for X-vines based on a random sample $\bm{X}_{1},\ldots,\bm{X}_{n}$ from $F$ involves:
@@
Selecting parametric (tail) copula families given a truncated regular vine sequence.
We first consider the specification of $\mathcal{R}$ in Definition 5.5 in tree $T_{1}$. Let $\mathcal{B}_{\mathcal{R},1:T}=(\mathcal{B}_{\mathcal{R},1},\ldots,\mathcal{B}_{\mathcal{R},T})$ be the list of candidate bivariate tail copula families. For each edge $e=(e_{a},e_{b})\in E_{1}$, specifying $\mathcal{R}$ involves choosing the bivariate tail copula family among $\mathcal{B}_{\mathcal{R},1:T}$. Similar to the idea of averaged maximum pseudo-likelihood estimates in Section 7.2, we use the averaged-AIC value for selecting bivariate tail copula families. For each edge $e$ and each $\mathcal{B}_{\mathcal{R},t}$, $t=1,\ldots,T$, we obtain two maximum pseudo-likelihood estimates $\hat{\theta}_{e}^{(t,a_{e})}$ and $\hat{\theta}_{e}^{(t,b_{e})}$, derived through the maximisation of log-likelihood functions on product spaces; see (7.3). The averaged-AIC value is $\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{R},t}}\right)=2\nu^{(t)}-\tfrac{1}{2}\left\{\log\mathcal{L}_{\mathcal{R},e}\left({\hat{\theta}_{e}^{(t,a_{e})}}\right)+\log\mathcal{L}_{\mathcal{R},e}\left({\hat{\theta}_{e}^{(t,b_{e})}}\right)\right\}$, where $\nu^{(t)}$ is the number of parameters in $\mathcal{B}_{\mathcal{R},t}$. We select the bivariate tail copula family with the lowest averaged-AIC among $\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{R},1}}\right),\ldots,\operatorname{AIC}_{e}\left({\mathcal{B}_{\mathcal{R},T}}\right)$.
Similarly, to specify $\mathcal{C}$ and select the bivariate parametric copula family among a list of candidates for each edge in $E_{2}\cup\cdots\cup E_{d-1}$, we follow common practice in vine copula modelling and choose the family with the lowest AIC; see Brechmann (2010) and Czado (2019, Section 8.1).
@@
Selecting the regular vine sequence.
@@
To select the first tree, $T_{1}$, on the node set $N_{1}=[d]$, we follow a procedure as in Engelke and Volgushev (2022) and Hu et al. (2024). For every pair $\left\{{a,b}\right\}$ of distinct elements in $[d]$, let $w_{a,b}$ be a nonnegative weight derived from the data. We use an empirical version of the tail dependence coefficient $\chi_{a,b}=R_{a,b}(1,1)$, setting $w_{a,b}$ equal to $\widehat{\chi}_{a,b}=\frac{1}{k}\sum_{i=1}^{n}\operatorname{\mathds{1}}\left\{{\widehat{U}_{i,a}\leqslant k/n,\widehat{U}_{i,b}\leqslant k/n}\right\}$, where $\widehat{U}_{i,a}$ and $\widehat{U}_{i,b}$ are defined as in (7.1). Another possible edge weight could be the empirical extremal variogram as in Engelke and Volgushev (2022). For subsequent trees $T_{2},\ldots,T_{d-1}$, the edge weight is chosen to be the absolute value of the empirical Kendall’s tau. All trees are selected as maximum spanning trees; see, for example, Czado (2019, Section 8.3).
Selecting truncated regular vine tree sequences.
The model selection procedure described above may set all pair-copulas to the independence copula in the subsequent trees from $T_{q+1}$ to $T_{d-1}$ for a truncation level $q\in\left\{{1,\ldots,d-2}\right\}$. In this case, the resulting model corresponds to the truncated X-vine model in Definition 5.10. Besides an information criterion, we consider two additional criteria for selecting the independence copula at an edge $e$: when the effective sample size $|K_{D_{e}}|=n_{D_{e}}$ falls below a certain low value, or when the absolute value of the empirical Kendall’s tau is close to zero. Sparsity induced by a smaller effective sample size, is more likely when the total sample size $n$ is relatively small with respect to the dimension $d$.
@@
$\operatorname{mBIC}(q)=\sum_{j=2}^{q}\sum_{e\in E_{j}}\left[{\operatorname{\mathds{1}}(\theta_{e}\neq 0)\left\{{\log{n_{D_{e}}}-2\log\left({\frac{\psi_{e}}{1-\psi_{e}}}\right)}\right\}-2\log\left\{\frac{\mathcal{L}_{\mathcal{C},e}\left({\hat{\theta}_{e}}\right)}{(1-\psi_{e})^{-1}}\right\}}\right]$ (7.5)
with $\mathcal{L}_{\mathcal{C},e}$ as in (D.1). The selected truncation level $q^{*}$ is the value of $q$ in $\left\{{1,\ldots,d-1}\right\}$ that minimises $\operatorname{mBIC}(q)$. In practice, we will set $\psi_{0}=0.9$ as in Nagler et al. (2019).
@@ 8.1 Parameter estimation
We first consider the $5$-dimensional X-vine model $r$ in Fig. LABEL:fig:5d_Xvine_a. The bivariate tail copula densities $r_{a_{e},b_{e}}$ in the first tree $T_{1}=(N_{1},E_{1})$ are chosen from the Hüsler–Reiss model, the negative logistic, logistic, and Dirichlet families (Section 4), while the bivariate copula densities $c_{a_{e},b_{e};D_{e}}$ in the subsequent trees $T_{j}=(N_{j},E_{j})$ for $j\in\left\{{2,3,4}\right\}$ are taken from the Clayton, Gumbel, and Gaussian copula families. Fig. LABEL:fig:5d_Xvine_a shows the values specified for the tail dependence coefficient $\chi=R(1,1)$ or for Kendall’s $\tau$ for each edge. The formulas connecting the parameters of the families of tail copula densities with $\chi$ are given in Appendix E.1 in the supplement.
@@
Figure 2: (a) An X-vine specification $(\mathcal{V},\mathcal{R},\mathcal{C})$ on $d=5$ variables. The bivariate tail copula densities $r_{a_{e},b_{e}}$ in the first tree $T_{1}$ are selected from the Hüsler–Reiss (HR), negative logistic (NL), logistic (L), and Dirichlet (Diri) parametric families, each with specified tail dependence coefficients $\chi$. The bivariate copula densities $c_{a_{e},b_{e};D_{e}}$ in the subsequent trees $T_{2},T_{3},T_{4}$ are chosen from the Clayton (Clay), Gumbel (Gum), and Gaussian (Ga) copula families, each with specified Kendall’s tau $\tau$. (b) Box-plots ($\blacksquare$ for $T_{1}$blacksquare for $T_{2}$blacksquare for $T_{3}$blacksquare for $T_{4}$) of dependence measure estimates from bivariate parametric families selected sequentially from the data for the X-vine specification in (a). The red lines indicate the specified parameter values.
Implementing simulation algorithms as detailed in Section 6 and Appendix A in the supplement, we generate multivariate inverted Pareto random samples $\bm{Z}_{1},\ldots,\bm{Z}_{n}$ associated with $r$. As in Eq. (7.1), we transform to $\widehat{\bm{U}}_{1},\ldots,\widehat{\bm{U}}_{n}$, with $\hat{U}_{i,j}$ based on the rank of $Z_{i,j}$ among $Z_{1,j},\ldots,Z_{n,j}$. We then take the sub-sample $\{\widehat{\bm{Z}}_{i}\}_{i\in K}$ with $K=\bigcup_{j\in[d]}K_{j}$ and $K_{j}$ as in (7.2).
@@ 8.2 Selecting parametric (tail) copula families
We assess algorithm effectiveness in selecting bivariate parametric families for each edge in each tree, using the X-vine specification in Fig. LABEL:fig:5d_Xvine_a. While the flexibility of X-vine models allows us to consider any bivariate parametric family, we simplify the process by considering a catalogue of four candidate tail copula models for $T_{1}$: the Hüsler–Reiss, logistic, negative logistic, and Dirichlet models, along with a catalogue of nine candidate pair-copula families for $T_{2},T_{3},T_{4}$: Independence, Gaussian, Clayton, Survival Clayton, Gumbel, Survival Gumbel, Frank, Joe, and Survival Joe copulas, as implemented in the R package VineCopula (Nagler et al., 2023).
@@
The lower proportions observed for $e_{23}$ and $e_{45}$ in $E_{1}$ result from the challenge in distinguishing between the logistic and negative logistic models. Overall, the proportion of accurately selected families decreases with declining effective sample size and Kendall’s tau estimate across tree levels. We observe the lowest proportion for the deepest edge $e_{15;234}$ where the corresponding pair-copula exhibits weak dependence. For this edge, the algorithm selects the independence copula 110 times out of 200.
We also investigate the effective sample sizes for each tree in Section F.1.4 of the supplement.
@@ 8.3 Tree selection and truncation
To explore truncated X-vine models, we use the mBIC in (7.5) with $\psi_{0}=0.9$ as in Nagler et al. (2019). The estimated mBIC-optimal truncation level is $q^{*}=19$ (dotted line in Fig. LABEL:fig:mBIC50dim). We assess the goodness of fit by comparing pairwise $\chi$-values from the true X-vine model with those from the fitted 50-dimensional X-vine model via Monte Carlo simulations (as explained in Section E.2 in the supplement) in Fig. LABEL:fig:Chi50dim. For the truncated X-vine model, the $\chi$-plot in Fig. LABEL:fig:Chi50Trunc resembles that of the full model but exhibits more variability, particularly for lower $\chi$-values.
Figure 3: Simulation study: (a) The mBIC plotted across tree levels with a dotted line indicating the selected mBIC-optimal truncation level of $q^{*}=19$. (b) $\chi$-plot comparing pairwise $\chi$ values from the true X-vine model to those obtained from the fitted 50-dimensional X-vine model via Monte Carlo simulations. (c) Similar to (b) but for the truncated fitted X-vine model.